OpenStudy (anonymous):

is the product of (a+b)(a-b) a perfect squared trinomal always sometimes oenever

OpenStudy (anonymous):

i need help plzz

OpenStudy (anonymous):

\[\left( a+b \right)\left( a-b \right)=a ^{2}-b ^{2}\]

OpenStudy (anonymous):

so is it always ?